# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework

### v2

@@ 1 Introduction
Accurate time series forecasting is critical in domains such as industrial monitoring (Jean-Pierre et al., 2024; Zhao et al., 2024b), energy systems (Yao et al., 2025; Nascimento et al., 2023; Hu, 2024), and healthcare (He & Chiang, 2025; Tang et al., 2024), where predictive insights directly influence operational safety, resource efficiency, and long-term outcomes. Yet, real-world time series present several challenges that are often oversimplified by current modeling approaches.
@@
As with interdependencies and sampling variability, missing data represents another key challenge in time series modeling. Real-world signals often contain long contiguous intervals of missing observations, often arising from maintenance issues, sensor malfunctions, or communication failures (You et al., 2025; Nejad et al., 2024). In such cases, naive interpolation, which is commonly applied during preprocessing, can be unreliable or misleading, particularly for dynamic signals. More robust approaches can instead benefit from leveraging cross-channel correlations to infer missing dynamics, rather than relying solely on extrapolation from partially observed inputs.
@@ 2 Related Work
Channel-Dependent Strategies for Multivariate Time Series
Channel-Dependent (CD) strategies jointly model all channels to capture inter-channel dependencies, and has been widely adopted across various neural architectures, including TCN/CNNs (Luo & Wang, 2024; Wu et al., 2023), MLPs (Han et al., 2024a; Huang et al., 2024), Transformers (Cheng et al., 2024; Liu et al., 2024b; Shu & Lampos, 2024; Wang et al., 2024e; c; Zhang & Yan, 2023; Liu et al., 2024c; Yang et al., 2024; Yu et al., 2023), and GNNs (Huang et al., 2023; Yi et al., 2023; Cao et al., 2020; Liu et al., 2022). Generative approaches such as COSCI-GAN (Seyfi et al., 2022) also employ CD principles, emphasizing cross-channel coherence in the context of sequence generation. These approaches are particularly effective when channels are correlated, since sparsely observed or low-resolution signals can benefit from denser ones through cross-channel interactions. In practice, however, most CD strategies only partially address real-world challenges: they typically assume aligned sampling across channels, are not designed to handle block-wise missing intervals and can be sensitive to noisy inter-dependencies. In contrast, our method utilizes channel tokens that explicitly encode cross-channel relationships while remaining robust under asynchronous sampling and block-wise missingness.
@@
Irregularly sampled multivariate time series are characterized by channels observed at non-uniform and often unpredictable time points, resulting in irregular gaps within individual channels. A number of recent works (Che et al., 2018; Shukla & Marlin, 2021; Li et al., 2023a; Chen et al., 2023; Zhang et al., 2023; 2024; Mercatali et al., 2024; Liu et al., 2024a; 2025; Kim & Lee, 2024; Klötergens et al., 2025) address this challenge by developing models that operate directly on non-uniform time intervals, often using time encodings, interpolation-based alignment, attention mechanisms, or graph-based structures. Models such as Raindrop (Zhang et al., 2022) and Hi-Patch (Luo et al., 2025) are tailored for highly sparse settings, focusing on point-wise inter-channel relations rather than capturing temporal patterns. As a result, they are less suited to structured multi-source asynchrony and do not align with our proposed practical setting, which unifies three core challenges in continuous, long-term forecasting. In contrast, we target a real-world setting, including fixed but distinct sampling periods across channels, as categorized in the recent Time-IMM dataset (Chang et al., 2025).
@@ 3 Problem Formulation
We consider a multivariate time series forecasting task where each channel $i\in\{1,\dots,N\}$ is sampled at a distinct, fixed period $s_{i}$. Let $s_{\min}=\min(s_{1},\dots,s_{N})$ denote the shortest sampling period, and define the sampling factor as $r_{i}=s_{i}/s_{\min}$. The input sequence for channel $i$ over a fixed observation window of length $L$ is represented as $\mathbf{x}^{(i)}=[x^{(i)}_{1},x^{(i)}_{2},\dots,x^{(i)}_{L_{i}}]\in\mathbb{R}^{L_{i}}$, where $L_{i}=\lfloor L/r_{i}\rfloor$. Given a prediction window of length $H$, the forecast horizon for channel $i$ becomes $\mathbf{y}^{(i)}=[y^{(i)}_{1},y^{(i)}_{2},\dots,y^{(i)}_{H_{i}}]\in\mathbb{R}^{H_{i}}$, where $H_{i}=\lfloor H/r_{i}\rfloor$. This implies that channels with longer sampling periods contain fewer observations within the same time window, resulting in shorter prediction horizons $H_{i}$. Note that this adjustment is made in the data point domain, not the time domain, so that the actual time horizon covered by each channel’s prediction remains consistent. We formulate a forecasting model $f_{\theta}$ parameterized by $\theta$ that predicts the future values by leveraging sampling period information from all channels: $\hat{\mathbf{y}}=f_{\theta}(\{(\mathbf{x}^{(i)},s_{i})\}_{i=1}^{N})$. Here, $s_{i}$ enables the model to account for the temporal resolution and data density of each input channel. The learning objective is to minimize the mean squared error across all channels, each of which may have a different prediction length $H_{i}$ due to its unique sampling period. The total training loss is computed by first calculating the average MSE over each channel’s prediction horizon, then averaging across channels. We refer to this as Channel-aggregated MSE (CMSE):
@@
In many real-world scenarios, observations exhibit persistent sparsity due to temporary failures, communication losses, or maintenance, leading to block-wise missing intervals at test time. This type of missingness is typically easy to detect, as most channels rarely show prolonged zero readings unless the device is disconnected or malfunctioning. Accordingly, we introduce a binary mask as auxiliary metadata to indicate whether each channel contains valid observations:
$\mathbf{m}^{(i)}\in\{0,1\}^{L_{i}},\quad m^{(i)}_{j}=\begin{cases}1,&\text{if }x_{j}^{(i)}\text{ is observed},\\ 0,&\text{if }x_{j}^{(i)}\text{ is missing}\end{cases}.$ (2)
The binary mask $\mathbf{m}^{(i)}$ indicates whether each patch in channel i should be masked or retained. Conventional patch-based methods cannot handle variable numbers of patches per channel, and thus are unable to leverage this binary mask. Instead, they typically fill missing entries within each patch using zeros or mean values. In contrast, our model adopts a non-overlapping patch representation that supports variable patch lengths and counts across channels. This enables the binary mask to be effectively used for removing fully unobserved patches from the input, rather than merely filling in missing values. Accordingly, the forecasting model is extended to accommodate masking-aware inputs across all channels: $\hat{\mathbf{y}}=f_{\theta}(\{(\mathbf{x}^{(i)},\mathbf{m}^{(i)},s_{i})\}_{i=1}^{N})$. This formulation supports forecasting under heterogeneous horizons from channel-wise asynchronous sampling, while also accommodating channel-specific missingness in the input at test time. The forecast horizon is assumed to be fully observed, since missing intervals in the future cannot be predicted in advance.
@@ 4 ChannelTokenFormer
Prior frequency-guided patching (e.g., Moirai (Woo et al., 2024), LightGTS (Wang et al., 2025)) in time series foundation models targeted variable-length segmentation for univariate streams with heterogeneous granularities. In contrast, we repurpose this frequency-guided granularity selection for a channel-wise asynchronous multivariate setting. This enables non-aligned, per-channel patching while sharing projection layers (i.e., tokenizers) across equal patch lengths and introducing dedicated channel tokens, thereby preserving efficiency. Specifically, each input channel $\mathbf{x}^{(i)}$ is adaptively patched by estimating its dominant frequency via the Fast Fourier Transform (FFT). If a clear peak is identified, the patch length is set to the corresponding period; otherwise, it is chosen through a sampling-aware fallback. Full details on thresholds, dataset-specific examples, and the fallback procedure are provided in Appendix A.2. This frequency-aware patching allows each channel to be divided into non-overlapping local patches that reflect its temporal characteristics and are subsequently summarized by channel tokens. Each local patch $\mathbf{p}^{(i)}_{j}\in\mathbb{R}^{\ell_{i}}$ is projected into a fixed-dimensional token $\mathbf{L}^{(i)}_{j}\in\mathbb{R}^{d}$ using a linear layer parameterized by the patch length: $\mathbf{L}^{(i)}_{j}=\mathbf{W}_{\text{patch}}^{(\ell_{i})}\mathbf{p}^{(i)}_{j}$. In parallel, multiple channel tokens $\mathbf{C}^{(i)}=[\mathbf{C}^{(i)}_{1};\dots;\mathbf{C}^{(i)}_{C}]\in\mathbb{R}^{C\times d}$ are initialized per channel to serve as global representations, where $C$ denotes the number of channel tokens assigned to each channel. We add a fixed positional embedding $\mathbf{e}_{\text{pos}}$ and a learnable channel embedding $\mathbf{e}_{\text{ch}}^{(i)}$ that encodes channel $i$’s identity to each local token, while the channel token receives only the channel embedding.
@@
Building on channel-wise dynamic patching, our framework supports variable input lengths via patch masking and explicitly targets block-wise missingness at test time. Unlike prior uses of random patch masking that primarily serve as a regularizer to curb overfitting, we repurpose it as a training-time proxy for realistic test-time incompleteness: inference-time inputs may be partially missing, so the model is trained to handle such missing blocks by design. We adopt random patch masking, as introduced in PatchDropout (Liu et al., 2023), applied to a subset of channel-wise patches at the input stage during training to simulate block-wise missingness at test time. As illustrated in Figure 2, if a missing segment spans an entire patch (as indicated by a binary mask) at test time, the corresponding local token is removed at the input stage, thereby being excluded from the attention computation. This prevents fully unobserved patches from contributing spurious signals, minimizing the risk of propagating invalid information. Despite temporal gaps induced by masking, temporal ordering and context are retained through fixed positional embeddings. This perspective shift improves resilience to test-time missing intervals and, as a side effect, acts as an implicit regularizer that mitigates overfitting in low-resource settings.
@@ A Unified Mask-Guided Attention with Channel Tokens for Real-World Challenges
To jointly address three practical challenges in multivariate time series, we design a unified, mask-guided self-attention mechanism that integrates local and global representations within a single attention operation. As illustrated in Figure 2, this design unifies both token types in one attention step. Each channel $i\in\{1,\dots,N\}$ contributes two types of tokens: (i) local tokens $\mathbf{L}^{(i)}\in\mathbb{R}^{L^{(i)}\times d}$ representing patch-level embeddings, and (ii) global channel tokens $\mathbf{C}^{(i)}\in\mathbb{R}^{C\times d}$ summarizing high-level contextual information, where $L^{(i)}$ and $C$ denote the number of local and channel tokens, respectively. The full token sequence is constructed as $\mathbf{X}=[\mathbf{L}^{(1)};\mathbf{C}^{(1)};\dots;\mathbf{L}^{(N)};\mathbf{C}^{(N)}]\in\mathbb{R}^{T\times d}$, with total token count $T=\sum_{i=1}^{N}(L^{(i)}+C)$. This unified sequence is passed through a masked multi-head self-attention layer with residual connection: $\mathbf{X}_{\text{out}}=\mathbf{X}+\text{Attention}(Q,K,V)$, where $Q=\mathbf{X}\mathbf{W}_{Q}$, $K=\mathbf{X}\mathbf{W}_{K}$, and $V=\mathbf{X}\mathbf{W}_{V}$ are linear projections of $\mathbf{X}$, and the attention is computed as $\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+\mathbf{M}\right)V$. Here, $\mathbf{M}\in\mathbb{R}^{T\times T}$ encodes structural constraints on how tokens can attend to one another, reflecting their types, channels, and masking. Our tailored masking scheme constructs $\mathbf{M}$ as illustrated in Figure 3: (1) Local tokens attend only to other local tokens within the same channel, enabling intra-temporal modeling. (2) Channel tokens attend to their own local tokens and to other channels’ tokens, but are not accessible to local tokens due to their read-only role in the attention mechanism. (3) Channel tokens do not attend to themselves; for each query $\mathbf{C}^{(i)}$, the key $\mathbf{C}^{(i)}$ is excluded to avoid self-reinforcement and to encourage informative cross-channel interaction. Our unified masking strategy, inspired by read-only prompt attention (Lee et al., 2023) and attention control techniques (Kim et al., 2025), is implemented without modifying the standard Transformer architecture. Stacked masked attention blocks with residual connections and layer normalization enable deeper modeling of temporal and cross-channel dependencies. Although channel-wise tokenization introduces additional parameters, the masking strategy ensures computation remains efficient by restricting attention to necessary token interactions, while scaling reliably to realistic channel counts and sequence lengths, as shown in Appendix C.3.
@@ 5 Experiments
To assess the effectiveness of ChannelTokenFormer under our proposed problem settings, we conduct extensive experiments on four widely used multivariate time series benchmarks adapted to practical conditions, and on two real-world datasets from air monitoring and the LNG cargo handling system. Additional results under conventional settings are provided in Appendix D.
@@
We compare CTF against representative state-of-the-art models across architectures and methods for irregularity and missingness. Transformer baselines include TimeXer (Wang et al., 2024e), iTransformer (Liu et al., 2024c), and PatchTST (Nie et al., 2023). We also cover CNNs, GNNs, and MLPs, including TimesNet (Wu et al., 2023), CrossGNN (Huang et al., 2023), TimeMixer++ (Wang et al., 2024a), and DLinear (Zeng et al., 2023). We further include irregular-sampling methods such as ContiFormer (Chen et al., 2023), t-PatchGNN (Zhang et al., 2024), and Hi-Patch (Luo et al., 2025) as well as a missingness-robust approach, BiTGraph (Chen et al., 2024b).
@@
Baseline models for regular time series assume uniformly sampled and fully observed inputs. To satisfy this requirement, we employ linear interpolation to recover values unobserved at timestamps not covered by channel-specific sampling periods. To isolate architectural effects from interpolation, we further implement interpolation-free variants, including a version of TimeXer (Wang et al., 2024e) modified to handle temporally non-aligned cross-channel inputs. Implementation details are provided in Appendix C.4. In contrast, irregular or missing-aware baselines do not require interpolation, as they operate solely on observed values. To simulate block-wise missingness at test time, we randomly mask contiguous input regions. For fairness, regular baselines without explicit missing-value handling are also provided with linearly interpolated values for these masked regions, consistent with their training procedure. For each dataset, input and prediction lengths are selected to reflect the channel-wise temporal resolution. Despite heterogeneous sampling periods across channels, it is practical to define input and prediction windows by a fixed time duration instead of requiring the same counts of sample points across channels. The actual number of input and output points per channel is determined by its sampling period, with shorter periods yielding more points over the same window.

### v4

@@ 1 Introduction
Accurate time series forecasting is critical in domains such as industrial monitoring (Jean-Pierre et al., 2024; Zhao et al., 2024b), energy systems (Yao et al., 2025; Nascimento et al., 2023; Hu, 2024), and healthcare (He & Chiang, 2025; Tang et al., 2024), where predictive insights directly influence operational safety, resource efficiency, and long-term outcomes. Yet, real-world time series forecasting faces many practical challenges, including irregular sampling, concept drift within channels, and uncertainty quantification. Among these, complex interdependencies among channels, channel-wise asynchrony, and block-wise missingness are often oversimplified by current modeling approaches, which typically assume channel-independent, regularly sampled, and complete time series. In this work, we explicitly address these three key challenges within a unified framework.
@@
As with interdependencies and sampling variability, missing data represents another key challenge in time series modeling. In practice, discrete missing values are already present in standard benchmarks such as ETTm, and modern neural time-series models can exhibit a certain degree of implicit robustness to such patterns (see Appendix C.6). However, real-world signals often contain long contiguous intervals of missing observations, often arising from maintenance issues, sensor malfunctions, or communication failures (You et al., 2025; Nejad et al., 2024). In such cases, naive interpolation, which is commonly applied during preprocessing, can be unreliable or misleading, particularly for dynamic signals. More robust approaches can instead benefit from leveraging cross-channel correlations to infer missing dynamics, rather than relying solely on extrapolation from partially observed inputs.
@@ 2 Related Work
Channel-Dependent (CD) Strategies for Multivariate Time Series
CD strategies jointly model all channels to capture inter-channel dependencies, and has been widely adopted across various neural architectures, including TCN/CNNs (Luo & Wang, 2024; Wu et al., 2023), MLPs (Han et al., 2024a; Huang et al., 2024), Transformers (Cheng et al., 2024; Liu et al., 2024b; Shu & Lampos, 2024; Wang et al., 2024e; c; Zhang & Yan, 2023; Liu et al., 2024c; Yang et al., 2024; Yu et al., 2023), and GNNs (Huang et al., 2023; Yi et al., 2023a; Cao et al., 2020; Liu et al., 2022). Some recent models (Chen et al., 2024a; Qiu et al., 2025) refine CD strategies by learning sparse or clustered inter-channel structures. Moreover, TimeFilter (Hu et al., 2025) constructs a spatial–temporal graph over patches and filters out spurious dependencies in a patch-specific manner. Generative approaches such as COSCI-GAN (Seyfi et al., 2022) also employ CD principles, emphasizing cross-channel coherence in the context of sequence generation. These approaches are particularly effective when channels are correlated, since sparsely observed or low-resolution signals can benefit from denser ones through cross-channel interactions. In practice, however, most CD strategies only partially address real-world challenges: they typically assume aligned sampling across channels, are not designed to handle block-wise missing intervals and can be sensitive to noisy inter-dependencies. In contrast, our method utilizes channel tokens that explicitly encode cross-channel relationships while remaining robust under asynchronous sampling and block-wise missingness.
@@
Irregularly sampled multivariate time series are characterized by channels observed at non-uniform and often unpredictable time points, resulting in irregular gaps within individual channels. A number of recent works (Che et al., 2018; Shukla & Marlin, 2021; Li et al., 2023a; Chen et al., 2023; Zhang et al., 2023; 2024; Mercatali et al., 2024; Liu et al., 2024a; 2025; Kim & Lee, 2024; Klötergens et al., 2025) address this challenge by developing models that operate directly on non-uniform time intervals, often using time encodings, interpolation-based alignment, attention mechanisms, or graph-based structures. Models such as Raindrop (Zhang et al., 2022b) and Hi-Patch (Luo et al., 2025) are tailored for highly sparse settings, focusing on point-wise inter-channel relations rather than capturing temporal patterns. As a result, they are less suited to structured multi-source asynchrony and do not align with our proposed practical setting, which unifies three core challenges in continuous, long-term forecasting. In contrast, we target a real-world setting, including fixed but distinct sampling periods across channels, as categorized in the recent Time-IMM dataset (Chang et al., 2025).
@@ 3 Problem Formulation
We consider a multivariate time series forecasting task with $N$ channels, where each channel $i\in\{1,\dots,N\}$ is sampled at a distinct fixed period. Let $\{s_{i}\}_{i=1}^{N}$ denote the relative sampling periods across channels, normalized such that $\min(s_{1},\dots,s_{N})=1$. Given an input length $L$, the number of valid sampled points available for channel $i$ becomes $L_{i}=\lfloor L/s_{i}\rfloor$. Similarly, for a prediction horizon of length $H$, the number of future points to be predicted for channel $i$ is $H_{i}=\lfloor H/s_{i}\rfloor$. This adjustment is made in the data-point domain, not the time domain, so that the actual time span covered by the prediction remains consistent across channels despite different temporal resolutions.
Our learning objective computes the forecasting error only at valid (observable) positions for each channel, which is standard in irregular forecasting settings (Luo et al., 2025; Zhang et al., 2024). We refer to this as Channel-aggregated MSE (CMSE):
@@
In many real-world cases, observations exhibit persistent sparsity due to temporary failures, communication losses, or maintenance, leading to block-wise missing intervals at test time, as shown in Figure 1. This type of missingness is typically easy to detect, as most channels rarely show prolonged zero readings unless the device is disconnected or malfunctioning.
Conventional patch-based methods cannot handle variable numbers of patches per channel, and thus typically fill missing entries within each patch using zeros or mean values. In contrast, our model adopts a non-overlapping patch representation that supports variable patch lengths and counts across channels. This enables the removing fully unobserved patches from the input, rather than merely filling in missing values. This formulation supports forecasting under heterogeneous horizons from channel-wise asynchronous sampling, while also accommodating channel-specific missingness in the input at test time. The forecast horizon is assumed to be fully observed, since missing intervals in the future cannot be predicted in advance.
@@ 4 ChannelTokenFormer
Prior frequency-guided patching (e.g., Moirai (Woo et al., 2024), LightGTS (Wang et al., 2025)) in time series foundation models targeted variable-length segmentation for univariate streams with heterogeneous granularities. In contrast, we repurpose this frequency-guided granularity selection for a channel-wise asynchronous multivariate setting. The detailed procedure is described in Algorithm 1. This enables non-aligned per-channel patching while preserving efficiency by sharing projection layers across channels with the same patch length and introducing dedicated channel tokens. For each channel, we estimate a dominant period via FFT to determine the patch length, and use a sampling-aware fallback when no clear peak exists. We then split each channel into non-overlapping patches and map each patch to a local token. For each channel, we also initialize a small set of channel tokens. Each local token receives a fixed positional embedding $\mathbf{e}_{\text{pos}}$ and a learnable channel embedding $\mathbf{e}_{\text{ch}}^{(i)}$, while the channel token receives only $\mathbf{e}_{\text{ch}}^{(i)}$. Full details on thresholds, dataset-specific examples, and the fallback procedure are provided in Appendix A.2.
@@
Building on channel-wise dynamic patching, our framework supports variable input lengths via patch masking and explicitly targets block-wise missingness at test time. Unlike prior uses of random patch masking that primarily serve as a regularizer to curb overfitting, we repurpose it as a training-time proxy for realistic test-time incompleteness: inference-time inputs may be partially missing, so the model is trained to handle such missing blocks by design. We adopt random patch masking, as introduced in PatchDropout (Liu et al., 2023), applied to a subset of channel-wise patches at the input stage during training to simulate block-wise missingness at test time. As illustrated in Figure 2, if a missing segment spans an entire patch at test time, the corresponding local token is removed at the input stage, thereby being excluded from the attention computation. This prevents fully unobserved patches from contributing spurious signals, minimizing the risk of propagating invalid information. Despite temporal gaps induced by masking, temporal ordering and context are retained through fixed positional embeddings. This perspective shift improves resilience to test-time missing intervals and, as a side effect, acts as an implicit regularizer that mitigates overfitting in low-resource settings. This procedure is formalized in Algorithm 1.
@@ A Unified Mask-Guided Attention with Channel Tokens for Real-World Challenges
To jointly address three practical challenges in multivariate time series, we design a unified, mask-guided self-attention mechanism that integrates local and global representations within a single attention operation. As illustrated in Figure 2, this design unifies both token types in one attention step. Each channel $i\in\{1,\dots,N\}$ contributes two types of tokens: (i) local tokens $\mathbf{T}^{(i)}\in\mathbb{R}^{\mathcal{T}^{(i)}\times d}$ representing patch-level embeddings, and (ii) global channel tokens $\mathbf{C}^{(i)}\in\mathbb{R}^{m\times d}$ summarizing high-level contextual information, where $\mathcal{T}^{(i)}$ and $C$ denote the number of local and channel tokens, respectively. The full token sequence is constructed as:
$\mathbf{X}=[\mathbf{T}^{(1)};\mathbf{C}^{(1)};\dots;\mathbf{T}^{(N)};\mathbf{C}^{(N)}]\in\mathbb{R}^{\mathcal{T}\times d},\;\text{where}\;\mathcal{T}=\sum_{i=1}^{N}(\mathcal{T}^{(i)}+m).$ (2)
This unified sequence is passed through a masked multi-head self-attention layer with residual connection:
$\mathbf{X}_{\text{out}}=\mathbf{X}+\text{Attention}(Q,K,V)=\mathbf{X}+\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+\mathbf{M}\right)V.$ (3)
Here, $\mathbf{M}\in\mathbb{R}^{\mathcal{T}\times\mathcal{T}}$ encodes structural constraints on how tokens can attend to one another, reflecting their types, channels, and masking. Our tailored masking scheme constructs $\mathbf{M}$ as illustrated in Figure 3: (1) Local tokens attend only to other local tokens within the same channel, enabling intra-temporal modeling. (2) Channel tokens attend to their own local tokens and to other channels’ tokens, but are not accessible to local tokens due to their read-only role in the attention mechanism. (3) Channel tokens do not attend to themselves; for each query $\mathbf{C}^{(i)}$, the key $\mathbf{C}^{(i)}$ is excluded to avoid self-reinforcement and to encourage informative cross-channel interaction. Our unified masking strategy, inspired by read-only prompt attention (Lee et al., 2023) and attention control techniques (Kim et al., 2025), is implemented without modifying the standard Transformer architecture. Stacked masked attention blocks with residual connections and layer normalization enable deeper modeling of temporal and cross-channel dependencies. Although channel-wise tokenization introduces additional parameters and our unified attention remains quadratic in the total number of tokens, the proposed masking scheme focuses attention on the most relevant token interactions, and empirically does not lead to prohibitive inference latency for realistic channel counts and sequence lengths (see Appendix C.3).
@@ 5 Experiments
To assess the effectiveness of ChannelTokenFormer (CTF) under our proposed problem settings, we conduct extensive experiments on four widely used multivariate time series benchmarks adapted to practical conditions, and on two real-world datasets from air monitoring and the LNG cargo handling system. Additional results under conventional settings are provided in Appendix D.
@@
We compare CTF against representative state-of-the-art models across architectures and modeling strategies that target inter-channel dependencies, irregular sampling, or missingness. Transformer baselines include TimeXer (Wang et al., 2024e), iTransformer (Liu et al., 2024c), and PatchTST (Nie et al., 2023). We also cover CNNs, GNNs, and MLPs, including TimesNet (Wu et al., 2023), CrossGNN (Huang et al., 2023), TimeMixer++ (Wang et al., 2024a), and DLinear (Zeng et al., 2023). We further include strong recent channel-dependent methods, DUET (Qiu et al., 2025) and TimeFilter (Hu et al., 2025); irregular-sampling methods such as t-PatchGNN (Zhang et al., 2024) and Hi-Patch (Luo et al., 2025); and a missingness-robust approach, BiTGraph (Chen et al., 2024b).
@@
Baseline models for regular time series assume uniformly sampled and fully observed inputs. To satisfy this requirement, we employ linear interpolation to recover values unobserved at timestamps not covered by channel-specific sampling periods. To isolate architectural effects from interpolation, we further implement interpolation-free variants, including a version of TimeXer (Wang et al., 2024e) modified to handle temporally non-aligned cross-channel inputs. Implementation details are provided in Appendix C.4. In contrast, irregular or missing-aware baselines do not require interpolation, as they operate solely on observed values. To simulate block-wise missingness at test time, we randomly mask contiguous input regions. For fairness, regular baselines without explicit missing-value handling are also provided with linearly interpolated values for these masked regions, consistent with their training procedure. For each dataset, input and prediction lengths are selected to reflect the channel-wise temporal resolution. Details for each dataset are in Appendix A.2. Despite heterogeneous sampling periods across channels, it is practical to define input and prediction windows by a fixed time duration instead of requiring the same counts of sample points across channels. The actual number of input and output points per channel is determined by its sampling period, with shorter periods yielding more points over the same window.